# Community detection in network using Szegedy quantum walk

### v1

@@ 4.5 Les Misérable graph
6(a) represents the “Les Misérable” graph. This network was demonstrated in [43] considering the characters in Victor Hugo’s novel “Les Misérable” as vertices. Any pair of characters that occur in the same chapter of the novel are connected by an edge. This graph has seventy seven vertices and two hundred fifty four edges.
Article [33] considers the Les Misérable graph for community detection. It determines six communities in the graph. We use Procedure 3 on the Les Misérable graph considering $q=\frac{1}{m}=0.0039$ for determining the communities. It generates eighteen communities in the graph. Among them there are three communities with vertices eight, twelve, and forty-four. Remaining fifteen communities have single vertices in each. In 6(d), we present the communities in the Les Misérable graph determined by our procedure. The coverage values corresponding to $q=0.0039$ is found to be $0.752$.
@@
(b) The graph in Subfigure 6(a) has a bar diagram that shows the edge probability in the limiting distribution $\Pi$.
(c) In 6(a), the path weight between various vertices of the graph is used to generate the communities.
(d) The three communities are represented by three distinct colors in the graph shown in Subfigure 6(a), while all other single-node clusters are shown by gray.
@@ 4.6 Dolphin’s social network
7(a) illustrates the dolphin’s social network which represents a social network among sixty two bottlenose dolphins, which are considered as nodes. Each of the one hundred fifty nine edges signifies a frequent social interaction between two dolphins spotted together more often than by chance [44].
Article [45] determines four communities in the dolphin’s social network. These communities consists of six, thirteen, twenty two and twenty one vertices. Considering $q=\frac{1}{m}=0.0063$ in Procedure 3, we obtain five distinct communities in dolphin social network. There are three communities having twelve, twenty-one, and twenty-seven vertices, and two communities consists of a single vertex in each. These single-vertex communities exist because the respective vertices are associated with several communities and hence cannot be unequivocally allocated to any one of the three communities. The coverage values corresponding to $q=0.0063$ is $0.824$. 7(a) presents the communities generated by our procedure.
@@
(b) The bar diagram for the graph in Subfigure 7(a) represents the edge probability in the limiting distribution $\Pi$.
(c) In 7(a), the path weight between various vertices of the graph is used to generate the communities.
(d) The three communities are represented by three distinct colors in the graph shown in Subfigure 7(a), while the other two single-node communities are shown by gray.
@@
4.7 Euroroad network
We consider Euroroad data from [46] for our experiment. The road interactions are represented by the vertices in the network. Every edge denotes a road that links two junctions. Our road network contains $1174$ nodes and $1417$ edges, which is not connected. Hence, we study the largest connected component in this network containing $1039$ vertices and $1305$ edges. This component is depicted in 8(a).
Applying Procedure 3, we find three hundred eighty-nine communities. The largest community consists of thirty-five vertices. The coverage values corresponding to $q=0.00077$ is found to be $0.540$.
(a) Connected Road network
(b) The bar diagram for the graph in Subfigure 8(a) represents the edge probability in the limiting distribution $\Pi$.
(c) In 8(a), the path weight between various vertices of the graph is used to generate the communities.
(d) All communities are represented by distinct colors in the graph shown in Subfigure 8(a), while the single-node communities are shown by gray.
Figure 8: Road Network are utilized to generate communities.
@@ 5 Conclusion
In this article, we introduce a new procedure that utilizes Szegedy quantum walks to detect the communities in a network. There are multiple models of quantum walks in the literature. We employ the Szegedy quantum walks on graphs for our investigation, as it is defined on all graphs.
The limiting probability distribution of the Szegedy quantum walks assists us in finding the communities. This distribution depends on the choice of the initial state of the quantum walker. We assume that the walker begins its journey from the vertices with the highest degree. Therefore, the initial quantum state as well as the limiting probability distribution depends on the structure of the graph. The limiting probability distribution provides the edge probabilities for each edge in the graph. We use it in Procedure 3 to extract the communities from the graph. This procedure is a combination of two subroutines, which are Procedure 1 and Procedure 2. In Procedure 1, we find the communities using quantum walks. Then, Procedure 2 refines the preliminary communities. We apply Procedure 3 on different graphs to identify the communities in them.
The problem of community detection is an important problem in the study of networks. The communities assists us to understand the structure of the networks. In the future, the interested reader may apply our community detection procedure in the investigation of different complex networks.
@@ Appendix A Dependency on the initial state
(a) An example graph
[…]
@@
(b) Limiting probability distributions on the graph in 9(a) for different choice of initial state.
Figure 9: We calculate limiting probability distributions mentioned in Definition 2 for a graph in 9(a) with different choice of initial states, which are mentioned in equation (5) and (15). We receive two different probability distributions in two different cases. They are in the table 9(b).
In this article, we consider a particular initial state, which is mentioned in equation (5). But, we had many other options of initial state of the walker. For example the equal superposition state as follows:
@@
Equation (15) indicates that the walker will start from all the edges selected uniformly.
@@
Appendix B Convergence of probability distributions
In Definition 2, we claim that the edge probability vectors, mentioned in Definition 1 will converge to a vector over time. The Markovian nature of the considered quantum walk is the key reason for this convergence. We present this incident via a number of diagrams. depicted in Figure 10.
We consider six network for our investigations. From every graph we select a random sample of four edges. For different time instance every edge has an edge probability value. We observe that all these probability values converges over time for any particular edge. Find Figure 10 for the detail descriptions.
@@ Appendix B Convergence of probability distributions
(g) Convergence of edge probability values over time to the limiting distribution $\Pi$ for a sample of edges of the road network, depicted in 8(a). We observe the convergence from $T=322$.

### v2

@@ 4.5 Les Misérable graph
6(a) represents the “Les Misérable” graph [49]. having $77$ vertices and $254$ edges.
Article [40] determines six communities in the graph. We use Procedure 3 considering $q=\frac{1}{m}=0.0039$ for determining the communities, that generates eighteen communities. Among them there are three communities with $8$, $12$, and $44$ vertices. Remaining $15$ communities have single vertices in each. In 6(c), we present these communities. The coverage values corresponding to $q=0.0039$ is $0.752$.
@@
(b) In 6(a), the path weight between various vertices of the graph is used to generate the communities.
(c) The three communities are represented by three distinct colors in the graph shown in Subfigure 6(a), while all other single-node clusters are shown by gray.
@@ 4.6 Dolphin’s social network
7(a) illustrates the dolphin’s social network [50]. Article [51] determines four communities in the network, that consists of $6$, $13$, $22$, and $21$ vertices.
Considering $q=\frac{1}{m}=0.0063$ in Procedure 3, we obtain $5$ communities. There are three communities having $12$, 2121, and $27$ vertices, and two communities consists of a single vertex in each. These single-vertex communities exist because the respective vertices are associated with several communities and hence cannot be unequivocally allocated to any one of the three communities. The coverage values corresponding to $q=0.0063$ is $0.824$. 7(a) presents the communities generated by our procedure.
@@
(b) In 7(a), the path weight between various vertices of the graph is used to generate the communities.
(c) The three communities are represented by three distinct colors in the graph shown in Subfigure 7(a), while the other two single-node communities are shown by gray.
@@ 5 Conclusion
In this article, we introduce a new procedure that utilizes Szegedy quantum walks to detect the communities in a network. There are multiple models of quantum walks in the literature. We employ the Szegedy quantum walks on graphs for our investigation.
The limiting probability distribution of the Szegedy quantum walks assists us in finding the communities. This distribution depends on the choice of the initial state of the quantum walker. We assume that the walker begins its journey from the vertices with the highest degree. The initial quantum state as well as the limiting probability distribution depends on the structure of the graph. We use it in Procedure 3 to extract the communities from the graph. This procedure is a combination of two subroutines, which are Procedure 1 and Procedure 2. In Procedure 1, we find the communities using quantum walks. Then, Procedure 2 refines the preliminary communities. We apply Procedure 3 on different graphs to identify the communities in them.
In the theory of complex network, the problem of community detection is an important problem, because the communities assists us to understand the structure of the networks. In the future, the interested reader may apply our community detection procedure for investigating the structure of different complex networks.
@@ Appendix A Dependency on the initial state
(a) A barbell graph.
Edge $\Pi^{\prime}$ $\Pi$
@@
(b) Limiting distributions.
Figure 8: A barbell graph and different limiting probability distributions for different initial states.
We consider an initial state, mentioned in equation (3). Consider another state
@@
The edge probability vectors, mentioned in Definition 1, converges to a limiting distribution, mentioned in Definition 2, which we can observe numerically. We consider a random sample of $4$ edges $\{(1,12),(23,24),(3,4),(6,7)\}$ from the karate club graph, which is depicted in 5(a). As none of them is incident to the highest degree vertices their probability values at $t=0$ is 0. Upto $t=50$ we present the changes in probability values which is depicted in Figure 9.